# One-loop contributions to $WWZ$ from new physics in the neutrino sector

### v1

@@ Abstract
The experimental confirmation of neutrino mass and mixing disagrees with the picture of the Standard Model, and thus incarnates a good motivation to look for unknown physics. A seesaw variant in which neutrino masses are radiatively generated, with neutrinos characterized by Majorana fields and where a set of three hypothetical neutral heavy leptons are present, has been considered in the present work. We have calculated, estimated, and analyzed the contributions from virtual light and heavy neutrinos to the gauge vertex $WWZ$, in the context of a future electron-positron collider. We have found that contributions to both $CP$-even and $CP$-odd anomalous couplings are generated. Our estimations indicate that contributions as large as ${}10^{-3}$ can be achieved in both cases. This is particularly interesting for the $CP$-odd effects, which are expected to appear in the framework of the Standard Model since the three-loop order, thus being quite suppressed. By considering the expected sensitivity of the International Linear Collider to this gauge vertex, we conclude that these new-physics effects would be within its reach for $e^{+}e^{-}$ collisions taking place at a center-of-mass energy of $800\,{\rm GeV}$.
@@ I Introduction
The evidence confirming neutrino oscillations, first achieved by experimental collaborations at the Super-Kamiokande experiment Kamiokande and at the Sudbury Neutrino Observatory SNO , showed that the neutrino sector of the Standard Model Glashow ; Salam ; Weinberg (SM) does not provide an accurate description of nature and thus requires to be extended. Massiveness of neutrinos, entailed by the phenomenon of neutrino oscillations Pontecorvo , might be introduced by assuming neutrinos to be characterized by Dirac fields Dirac and then endowing them with masses just the way it is done with all other fermions in the Yukawa sector of the SM. However, aiming at a more sort of natural definition of neutrino masses, in the sense of explaining their conspicuous tininess, alternative mechanisms have to be searched for. A well-known neutrino-mass generating mechanism is the seesaw MoSe1 ; MoSe2 , which propounds that neutrinos, assumed to be described by Majorana fermion fields Majorana , have such small masses due to some energy scale, $\Lambda$, associated to a yet-unknown high-energy physical description. Though, from a purely theoretical viewpoint, this explanation is quite appealing, it comes along with a practical drawback: the very large size of the high-energy scale $\Lambda$, imposed by current experimental bounds on neutrino masses KATRIN ; SDSSexperiment ; Planckexperiment , avoids any possibility of measuring the new-physics (NP) effects by current and future experimental facilities, perhaps even in the long term. In particular, the seesaw mechanism introduces a new set of heavy neutral leptons, here referred to as “heavy neutrinos”, with masses linked to those of the usual light neutrinos by $\Lambda$. In this context, the masses of such heavy neutrinos turn out to be gigantic, so producing them or measuring their quantum effects on SM observables does not seem to be achievable. Pursuing a seesaw-like neutrino-mass origin in which heavy neutrinos bear masses with a more reasonable size, seesaw variants have been conceived  [^1] For a review on seesaw variants, see Ref. CHLR and references therein.. Among the available seesaw-type mechanisms, we have considered the model given in Ref. Pilaftsis as the framework for the present investigation. The author of this reference introduced, in the context of the seesaw mechanism, a condition to render all tree-level light-neutrino mass-terms zero, while keeping masses of heavy neutrinos untouched. This breaks the strong link among heavy- and light-neutrino masses, which allows for quite smaller masses for the heavy neutral leptons. Then, masses of light neutrinos are rather generated by radiative corrections, which defines a new connection of neutrino masses, namely, in order for light-neutrinos to have adequate tiny masses, the set of heavy-neutrino masses has to be quasi-degenerate.
@@
where $\Gamma_{\sigma\rho\mu}^{\rm even}$ is associated to $CP$-conserving effects, whereas $\Gamma_{\sigma\rho\mu}^{\rm odd}$ characterizes physics not preserving $CP$ symmetry. If the $W$-boson external lines are taken on shell, while assuming that the external $Z$-boson is off shell, the $CP$-even and $CP$-odd terms, $\Gamma_{\sigma\rho\mu}^{\rm even}$ and $\Gamma_{\sigma\rho\mu}^{\rm odd}$, are respectively given by BGL ; HPZH ; BaZe
@@
Our interest, in the present investigation, focuses on the NP contributions from virtual Majorana neutrinos to $WWZ$, produced at one loop, for which we consider both light and heavy neutrinos, in the framework defined by the model of Ref. Pilaftsis . The associated NP contributions to the AC $\Delta\kappa$, $\Delta Q$, $\Delta\tilde{\kappa}$, and $\Delta\tilde{Q}$, characterizing this gauge vertex, are computed, estimated and analyzed. While both $CP$-conserving contributions, $\Delta Q$ and $\Delta\kappa$, are generated, the $CP$-odd $\Delta\tilde{Q}$ is found to vanish, so the only nonzero $CP$-violating contribution turns out to be $\Delta\tilde{\kappa}$. Anticipating the presumable presence of ultraviolet divergences, we utilize the dimensional-regularization method BoGi ; tHooVe . At the end of the day, we arrive at the conclusion that all the generated contributions to $\Delta Q$, $\Delta\kappa$, and $\Delta\tilde{\kappa}$ are ultraviolet finite. Feynman rules for Dirac fields differ from those for Majorana fermions DEHK ; GlZr . In particular, given some physical process, the Majorana framework usually involves a larger number of contributing diagrams than the Dirac case. In the present work, the set of Feynman diagrams contributing via virtual Majorana fermions includes (1) those diagrams which would contribute to the same process if such fermions were assumed to be of Dirac type, and (2) a set of extra diagrams to be taken into account. We call the extra diagrams, characteristic of Majorana fermions, “Majorana-type diagrams”, whereas those which occur in both the Dirac and Majorana cases are referred to as “Dirac-type diagrams”. We find that the contributions from Majorana-type diagrams exactly match those of Dirac-type diagrams. Our numerical analysis of the $WWZ$ vertex is carried out in the context of some future electron-positron colliding facility, such as the highly anticipated International Linear Collider (ILC) ILC1 ; ILC2 . We consider the process $e^{+}e^{-}\to W^{+}W^{-}$, which receives contributions from an $s$-channel diagram involving a virtual $Z$-boson line. Therefore, our calculation is performed under the assumption that the external $Z$-boson line is off the mass shell, in which case, in accordance with the notation defined by Fig. 1, the $WWZ$ vertex function $\Gamma^{WWZ}_{\sigma\rho\mu}$ depends on the squared 4-momentum $(2q)^{2}$. Moreover, the center-of-mass energy (CME), $\sqrt{s}$, of the corresponding $e^{+}e^{-}$ collision relates to this momentum as $s=4q^{2}$, so all contributions to AC are CME dependent. The present investigation aims at the calculation of such off-shell contributions, as well as its analysis and estimation, in the light of the expected ILC sensitivity ILC2 . Our estimations indicate that $CP$-preserving contributions $\Delta\kappa$ as large as $\sim 10^{-3}$ can be generated for a electron-positron collision at a CME of $800\,{\rm GeV}$, thus falling into the sensitivity region which is expected for the ILC. On the other hand, the only nonzero $CP$-odd contribution, $\Delta\tilde{\kappa}$, also reaches $\sim 10^{-3}$.

### v2

# One-loop contributions to $WWZ$ from a seesaw variant with radiatively-induced light-neutrino masses

@@ Abstract
The experimental confirmation of neutrino mass and mixing disagrees with the picture of the Standard Model, and thus provides a good motivation to look for unknown physics. A seesaw variant in which neutrino masses are radiatively generated, with neutrinos characterized by Majorana fields and where a set of three hypothetical neutral heavy leptons are present, has been considered in the present work. We have calculated, estimated, and analyzed the contributions from virtual light and heavy neutrinos to the gauge vertex $WWZ$, in the context of a future electron-positron collider. We have found that contributions to both $CP$-even and $CP$-odd anomalous couplings are generated. Our estimations indicate that contributions as large as ${}10^{-3}$ can be achieved in both cases. This is particularly interesting for the $CP$-odd effects, which are expected to appear in the framework of the Standard Model since the three-loop order, thus being quite suppressed. By considering the expected sensitivity of the International Linear Collider to this gauge vertex, we conclude that these new-physics effects would be within its reach for $e^{+}e^{-}$ collisions taking place at a center-of-mass energy of $800\,{\rm GeV}$.
@@ I Introduction
The evidence confirming neutrino oscillations, first achieved by experimental collaborations at the Super-Kamiokande experiment Kamiokande and at the Sudbury Neutrino Observatory SNO , showed that the neutrino sector of the Standard Model Glashow ; Salam ; Weinberg (SM) does not provide an accurate description of nature and thus requires to be extended. Massiveness of neutrinos, entailed by the phenomenon of neutrino oscillations Pontecorvo , might be introduced by assuming neutrinos to be characterized by Dirac fields Dirac and then endowing them with masses just the way it is done with all other fermions in the Yukawa sector of the SM. However, aiming at a more sort of natural definition of neutrino masses, in the sense of explaining their conspicuous tininess, alternative mechanisms have to be searched for. A well-known neutrino-mass generating mechanism is the seesaw MoSe1 ; MoSe2 ; PaSa , which propounds that neutrinos, assumed to be described by Majorana fermion fields Majorana , have such small masses due to some high-energy scale, $\Lambda$, associated to a yet-unknown physical description, beyond the Standard Model. The Weinberg operator Weinbergoperator , an effective-Lagrangian term DGMP with units $\big{(}{\rm mass}\big{)}^{5}$ and which introduces violation of lepton number, generates, driven by the Brout-Englert-Higgs mechanism EnBr ; Higgs , Majorana-mass terms for neutrinos, with masses given by $m_{\nu_{j}}\sim\frac{v^{2}}{\Lambda}$, where $v=246\,{\rm GeV}$ is the SM Higgs vacuum expectation value. This mass profile matches the one emerged from the seesaw mechanism, which, in addition, comes along with a set of heavy-neutral leptons, dubbed “heavy-neutrinos”, with masses $m_{N_{j}}\sim\Lambda$. In this context, current upper bounds on neutrino mass KATRIN ; SDSSexperiment ; Planckexperiment , lying within the sub-eV regime, push the energy scale $\Lambda$ towards enormous values, thus yielding huge heavy-neutrino masses. Therefore, direct production of heavy neutrinos or measurement of their quantum effects on SM observables do not seem to be achievable. So, while from a purely theoretical viewpoint, the seesaw explanation is quite appealing, it comes along with a practical drawback: the very large size of the high-energy scale $\Lambda$ avoids any possibility of measuring the new physics (NP) through current and future experimental facilities, perhaps even in the long term. Pursuing a seesaw-like neutrino-mass origin in which heavy neutrinos bear masses with a more reasonable size, seesaw variants have been conceived [^1] For a review on seesaw variants, see Ref. CHLR and references therein.. The inverse seesaw MoVa ; GoVa ; DeVa and the linear seesaw ALSV1 ; ALSV2 neutrino-mass-generating mechanisms are well-known instances. Among the available seesaw-type mechanisms, we have considered the model given in Ref. Pilaftsis as the framework for the present investigation. The author of this reference introduced, in the context of the seesaw mechanism, a condition to render all tree-level light-neutrino mass-terms zero, while keeping masses of heavy neutrinos untouched. This breaks the seesaw-type link among heavy- and light-neutrino masses, which allows for quite smaller masses for the heavy neutral leptons. Then, masses of light neutrinos are rather generated by radiative corrections, which defines a new connection of neutrino masses, namely, in order for light-neutrinos to have adequate tiny masses, the set of heavy-neutrino masses has to be quasi-degenerate.
@@
where $\Gamma_{\sigma\rho\mu}^{\rm even}$ is associated to $CP$-conserving effects, whereas $\Gamma_{\sigma\rho\mu}^{\rm odd}$ characterizes physics not preserving $CP$ symmetry. If the $W$-boson external lines are taken on shell, while assuming that the external $Z$-boson is off shell, the $CP$-even and $CP$-odd terms, $\Gamma_{\sigma\rho\mu}^{\rm even}$ and $\Gamma_{\sigma\rho\mu}^{\rm odd}$, are respectively given by [^2] A brief discussion on the effective-Lagrangian origin of both vertex functions $\Gamma^{\rm even}_{\sigma\rho\mu}$ and $\Gamma^{\rm odd}_{\sigma\rho\mu}$ is provided in Appendix A BGL ; HPZH ; BaZe
@@
Our interest, in the present investigation, focuses on the NP contributions from virtual Majorana neutrinos to $WWZ$, produced at one loop, for which we consider both light and heavy neutrinos, in the framework defined by the model of Ref. Pilaftsis . The associated NP contributions to the AC $\Delta\kappa$, $\Delta Q$, $\Delta\tilde{\kappa}$, and $\Delta\tilde{Q}$, characterizing this gauge vertex, are computed, estimated and analyzed. While both $CP$-conserving contributions, $\Delta Q$ and $\Delta\kappa$, are generated, the $CP$-odd $\Delta\tilde{Q}$ is found to vanish, so the only nonzero $CP$-violating contribution turns out to be $\Delta\tilde{\kappa}$. Anticipating the presumable presence of ultraviolet divergences, we utilize the dimensional-regularization method BoGi ; tHooVe . At the end of the day, we arrive at the conclusion that all the generated contributions to $\Delta Q$, $\Delta\kappa$, and $\Delta\tilde{\kappa}$ are ultraviolet finite. Feynman rules for Dirac fields differ from those for Majorana fermions DEHK ; GlZr . In particular, given some physical process, the Majorana framework usually involves a larger number of contributing diagrams than the Dirac case. In the present work, the set of Feynman diagrams contributing via virtual Majorana fermions includes (1) those diagrams which would contribute to the same process if such fermions were assumed to be of Dirac type, and (2) a set of extra diagrams to be taken into account. We call the extra diagrams, characteristic of Majorana fermions, “Majorana-type diagrams”, whereas those which occur in both the Dirac and Majorana cases are referred to as “Dirac-type diagrams”. We find that the contributions from Majorana-type diagrams exactly match those of Dirac-type diagrams. Our numerical analysis of the $WWZ$ vertex is carried out in the context of some future electron-positron colliding facility, such as the highly anticipated International Linear Collider (ILC) ILC1 ; ILC2 . We consider the process $e^{+}e^{-}\to W^{+}W^{-}$, which receives contributions from an $s$-channel diagram involving a virtual $Z$-boson line. Therefore, our calculation is performed under the assumption that the external $Z$-boson line is off the mass shell, in which case, in accordance with the notation defined by Fig. 1, the $WWZ$ vertex function $\Gamma^{WWZ}_{\sigma\rho\mu}$ depends on the squared 4-momentum $(2q)^{2}$. Moreover, the center-of-mass energy (CME), $\sqrt{s}$, of the corresponding $e^{+}e^{-}$ collision relates to this momentum as $s=4q^{2}$, so all contributions to AC are CME dependent. The present investigation aims at the calculation of such off-shell contributions, as well as its analysis and estimation, in the light of the expected ILC sensitivity ILC2 . Our estimations indicate that $CP$-preserving contributions $\Delta\kappa$ as large as $\sim 10^{-3}$ can be generated for an electron-positron collision at a CME of $800\,{\rm GeV}$, thus falling into the sensitivity region expected for the ILC. On the other hand, the only nonzero $CP$-odd contribution, $\Delta\tilde{\kappa}$, also reaches $\sim 10^{-3}$, which lies about one order of magnitude out of the reach of ILC expected sensitivity.